OpenStudy (anonymous):

someone help! What is the equation of the line that passes through (4, 2) and is parallel to 3x – 2y = –6?

OpenStudy (trojanpoem):

I wish I hadn't mistaken in anything. the slope of the line = -x / y = -3 / -2 = 3/2 Assume \[(a, b) \in line\] \[\frac{ b - 2 }{ a - 4 } = \frac{ 3 }{ 2 }\] 3a - 12 = 2b - 4 3 x - 2 y - 8 = 0